OpenStudy (anonymous):

Please Help for medal: Solve: x^2-3x-10=0

OpenStudy (anonymous):

"Factor the equation to (x - 5)(x + 2) = 0 Set both factors equal to 0 x - 5 = 0 x + 2 = 0 Solve for x and the answer is x = 5 x = -2" Does that help?

OpenStudy (anonymous):

\[x ^{2} - 3x - 10 = 0\] Factor each out into two paranthetical sub groups (x - 5)(x + 2) = 0 Set both groups = to 0. Solve for x

OpenStudy (anonymous):

\[\huge x^2-3x-10=0 \rightarrow x^2-5x+2x-10=0\] \[\huge \rightarrow x(x-5)+2(x-5)=0\] \[\huge \rightarrow (x+2)(x-5)=0\] \[\huge \rightarrow (x+2)=0 ---or ---(x-5)=0\] \[\huge \rightarrow x=-2 ---or ---x=5\] Hence x= -2, 5 are the required solutions